OpenStudy (anonymous):

Solve 9b^2 – 14b – 8 = 0

OpenStudy (anonymous):

TheViper are u gonna answer?

OpenStudy (theviper):

sorry i was not there

OpenStudy (anonymous):

can you answer it tho?

OpenStudy (anonymous):

9b^2 – 14b – 8 = 0 9b^2 – 18b+4b – 8 = 0 9b(b-2)+4(b-2) =0 (9b+4)(b-2)=0 therefore 9b+4 =0 and b = -4/9 or b-2 = 0 and b=2 solution (-4/9, 2)